Ethanol (C2H5OH) melts at -114 ∘C and boils at 78 ∘C. The enthalpy of fusion of ethanol is 5.02 kJ/mol, and its enthalpy of vaporization is 38.56 kJ/mol. The specific heat of solid and liquid ethanol are 0.97 J/g⋅K are 2.3 J/g⋅K respectively.

How much heat is required to convert 32.0 g of ethanol at -161 ∘C to the vapor phase at 78 ∘C?

Express your answer using two significant figures.
Answer #1

Ans :

Q1 = heat required to raise the temperature of solid ethanol to -114 C

= mass x specific heat x change in temperature

= 32.0 x 0.97 x ( -114 -(-161))

= 1458.88 J

Q2 = heat of fusion = mol x hfus

mol = mass / molar mass

= 32.0 / 46.07 = 0.6946 mol

Q2 = 0.6946 x 5.02

= 3.487 KJ

Q3 = heat required to raise the temperature of liquid ethanol to 78 C

= mass x specific heat x change in temperature

= 32.0 x 2.3 x ( 78 -(-114))

= 14131.2 J

Q4 = heat required to vaporise = mol x hvap

= 0.6946 x 38.56

= 26.78 KJ

So total heat = Q1 + Q2 + Q3 + Q4

= 1458.88 J + 3.487 KJ + 14131.2 J + 26.78 KJ

= 46 KJ